OpenStudy (anonymous):

a car travelling at 7.0m/s accelerates uniforly at 2.5m/s² to reach a speed of 12.0m/s. how long does it take for this acceleration to occur?

OpenStudy (anonymous):

t=(v-u)/a =(12-7)/2.5 = 5/2.5 = 2 s